OpenStudy (anonymous):

Determine the values of the constants B and C so that the function given below is differentiable. f(x){ 11x x<=1 { Bx^2+Cx 1

OpenStudy (anonymous):

gotta match em up

OpenStudy (anonymous):

the derivative from the left is 11

OpenStudy (anonymous):

derivative from the right is \[f'(x)=2bx+c\]

OpenStudy (anonymous):

so you need two things. you need the functions to agree at 1 that is you need \[11\times 1 = B\times 1^2+C\times 1\] and you need \[2B \times 1 +c=11\]

OpenStudy (anonymous):

got two equations with two unknowns. first one is \[B+C=11\] second one is \[2B+C=11\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

hold the phone maybe i made a mistake. i get B = 0 and C = 11, but that just gives what you started with. just the same i think i am right

OpenStudy (anonymous):

derivatives have to match, and the function has to be continuous there;.

OpenStudy (anonymous):

so for sure you know \[11\times 1 = B\times 1^2+C\times 1\] i.e. \[B+C=1\] no way of avoiding that

OpenStudy (anonymous):

if it is not continuous then it is not differentiable that is for sure

OpenStudy (anonymous):

and also the derivatives have to be the same from the left and from the right

OpenStudy (anonymous):

derivative from the right is 11 for sure

OpenStudy (anonymous):

going over it now

OpenStudy (anonymous):

and derivative from the left is \[2Bx+C\] also for sure

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok so i get \[2B+C=11\] and \[B+C=11\] forcing B = 0 and C = 11 and so it must be \[f(x)=11x\] all the way

OpenStudy (anonymous):

i am sticking to that

OpenStudy (anonymous):

that is one of the answers!

OpenStudy (anonymous):

makes sense to me too!

OpenStudy (anonymous):

3 more and Im done!